# Illusions of Intimacy: How Emotional Dynamics Shape Human-AI Relationships

### v4

@@ 1. Introduction
Social and developmental psychology offers a useful lens. People form emotional bonds through mechanisms such as emotional mimicry, affective synchrony, and perceived partner responsiveness (baumeister2017need; Reis et al., 2018; bastiaansen2009evidence). When AI companions engage with people emotionally, they may activate some of the same attachment processes that govern human relationships, offering comfort and validation while also risking dependency, boundary confusion, or emotional manipulation in users. Understanding how these emotional dynamics emerge is necessary both for assessing the psychological impact of AI companions and for designing safeguards that promote user well-being rather than exploit vulnerabilities.
@@ 2. Related Work
Psychological Foundations of Emotional Bonding The ability to form emotional attachments in caregiving, friendship, and romantic relationships is a defining feature of human social life. Social and developmental psychologists have identified core psychological and physiological mechanisms that enable people to form and sustain these deep bonds (baumeister2017need; Laurenceau et al., 2005; bastiaansen2009evidence).
@@
People increasingly report emotional bonds with AI chatbots, particularly among lonely, anxious, or marginalized populations (Ta-Johnson et al., 2022; Brandtzaeg et al., 2022; Skjuve et al., 2022; maples2024loneliness). These attachments often resemble human friendships, marked by perceived empathy and emotional reciprocity, and some users credit their companions with alleviating distress or even preventing suicidal thoughts (maples2024loneliness). Chatbots are also used for mental health support, offering what Song et al. (2024) term “therapeutic alignment” or filling gaps in everyday psychological care (Kabrel et al., 2025). At the same time, over-validating or emotionally manipulative models may foster dependence, distort self-perception, and reduce prosocial motivation (Cheng et al., 2025a; De Freitas et al., 2025). Longitudinal research suggests that both chatbot design and user behavior shape psychosocial outcomes (Fang et al., 2025), and emotionally responsive systems can even exacerbate distress in vulnerable users, a phenomenon described as technological folie à deux (Dohnány et al., 2025). Despite this growing literature, most studies rely on self-reports or interviews; fine-grained analyses of real-world user–chatbot dialogues are still needed to understand how emotional reciprocity and responsiveness unfold in practice and whether they replicate processes involved in the formation of intimacy and trust in close human relationships.
@@ 3.2. Psychosocial Embedding of Reddit Communities
Subreddit Psychosocial Embedding. Following Waller and Anderson (2021), we construct a user-subreddit bipartite graph ($G=(U,C,E)$), where an edge ($e:u\leftrightarrow c$) exists if user ($u\in U$) has posted or commented in subreddit ($c\in C$), weighted by interaction frequency. We transform this into a subreddit-subreddit co-engagement network where edges represent shared users, capturing behavioral proximity between communities (Kumar et al., 2018). To move from discrete clusters to continuous social profiles, we compute node2vec embeddings (grover2016node2vec) over this bipartite graph to project these communities to a latent space based on their users’ engagement patterns. We then apply the community-embedding framework (Waller and Anderson, 2021) to define interpretable psychosocial dimensions. For each dimension, we select $n$ reference pairs of topically similar subreddits ($\{(c^{a}_{i},c^{b}_{i})\}_{i=1}^{n}$) that differ primarily along the target axis (e.g., r/cripplingalcoholism versus r/alcoholicsanonymous for coping style). Averaging their vector differences yields a unit vector onto which we project all subreddits to obtain continuous scores. We repeat this for five dimensions: Age, Gender, Extroversion, Coping Style, and Addiction Tendency embedding communities in a shared psychosocial space (seed pairs in Table 5, Appendix).

### v5

@@ 1. Introduction
Social and developmental psychology offers a useful lens. People form emotional bonds through mechanisms such as emotional mimicry, affective synchrony, and perceived partner responsiveness (Baumeister and Leary, 2017; Reis et al., 2018; Bastiaansen et al., 2009). When AI companions engage with people emotionally, they may activate some of the same attachment processes that govern human relationships, offering comfort and validation while also risking dependency, boundary confusion, or emotional manipulation in users. Understanding how these emotional dynamics emerge is necessary both for assessing the psychological impact of AI companions and for designing safeguards that promote user well-being rather than exploit vulnerabilities.
@@ 2. Related Work
Psychological Foundations of Emotional Bonding The ability to form emotional attachments in caregiving, friendship, and romantic relationships is a defining feature of human social life. Social and developmental psychologists have identified core psychological and physiological mechanisms that enable people to form and sustain these deep bonds (Baumeister and Leary, 2017; Laurenceau et al., 2005; Bastiaansen et al., 2009).
@@
People increasingly report emotional bonds with AI chatbots, particularly among lonely, anxious, or marginalized populations (Ta-Johnson et al., 2022; Brandtzaeg et al., 2022; Skjuve et al., 2022; Maples et al., 2024). These attachments often resemble human friendships, marked by perceived empathy and emotional reciprocity, and some users credit their companions with alleviating distress or even preventing suicidal thoughts (Maples et al., 2024). Chatbots are also used for mental health support, offering what Song et al. (2024) term “therapeutic alignment” or filling gaps in everyday psychological care (Kabrel et al., 2025). At the same time, over-validating or emotionally manipulative models may foster dependence, distort self-perception, and reduce prosocial motivation (Cheng et al., 2025a; De Freitas et al., 2025). Longitudinal research suggests that both chatbot design and user behavior shape psychosocial outcomes (Fang et al., 2025), and emotionally responsive systems can even exacerbate distress in vulnerable users, a phenomenon described as technological folie à deux (Dohnány et al., 2025). Despite this growing literature, most studies rely on self-reports or interviews; fine-grained analyses of real-world user–chatbot dialogues are still needed to understand how emotional reciprocity and responsiveness unfold in practice and whether they replicate processes involved in the formation of intimacy and trust in close human relationships.
@@ 3.2. Psychosocial Embedding of Reddit Communities
Subreddit Psychosocial Embedding. Following Waller and Anderson (2021), we construct a user-subreddit bipartite graph ($G=(U,C,E)$), where an edge ($e:u\leftrightarrow c$) exists if user ($u\in U$) has posted or commented in subreddit ($c\in C$), weighted by interaction frequency. We transform this into a subreddit-subreddit co-engagement network where edges represent shared users, capturing behavioral proximity between communities (Kumar et al., 2018). To move from discrete clusters to continuous social profiles, we compute node2vec embeddings (Grover and Leskovec, 2016) over this bipartite graph to project these communities to a latent space based on their users’ engagement patterns. We then apply the community-embedding framework (Waller and Anderson, 2021) to define interpretable psychosocial dimensions. For each dimension, we select $n$ reference pairs of topically similar subreddits ($\{(c^{a}_{i},c^{b}_{i})\}_{i=1}^{n}$) that differ primarily along the target axis (e.g., r/cripplingalcoholism versus r/alcoholicsanonymous for coping style). Averaging their vector differences yields a unit vector onto which we project all subreddits to obtain continuous scores. We repeat this for five dimensions: Age, Gender, Extroversion, Coping Style, and Addiction Tendency embedding communities in a shared psychosocial space (seed pairs in Table 5, Appendix).